# Distributed Perceptron under Bounded Staleness, Partial Participation, and Noisy Communication

### v1

@@ Abstract
We study a semi-asynchronous client-server perceptron trained via iterative parameter mixing (IPM-style averaging): clients run local perceptron updates and a server forms a global model by aggregating the updates that arrive in each communication round. The setting captures three system effects in federated and distributed deployments: (i) stale updates due to delayed model delivery and delayed application of client computations (two-sided version lag), (ii) partial participation (intermittent client availability), and (iii) imperfect communication on both downlink and uplink, modeled as effective zero-mean additive noise with bounded second moment. We introduce a server-side aggregation rule called staleness-bucket aggregation with padding that deterministically enforces a prescribed staleness profile over update ages without assuming any stochastic model for delays or participation. Under margin separability and bounded data radius, we prove a finite-horizon expected bound on the cumulative weighted number of perceptron mistakes over a given number of server rounds: the impact of delay appears only through the mean enforced staleness, whereas communication noise contributes an additional term that grows on the order of the square root of the horizon with the total noise energy. In the noiseless case, we show how a finite expected mistake budget yields an explicit finite-round stabilization bound under a mild fresh-participation condition.
@@ I Introduction
System effects: staleness, partial participation, and noisy links
Practical federated and distributed deployments must contend with three persistent phenomena: (i) staleness (version lag), as in bounded-staleness frameworks such as SSP [6, 7] and online learning with delayed feedback [14, 15]; (ii) partial participation, where only a subset of clients return updates each round [5, 11, 12]; and (iii) communication noise on downlink/uplink, especially in wireless settings [8, 9, 10].
Approach and contributions
@@
Notation
@@
II System Model and Staleness-Profile IPM Perceptron
II-A Data and separability
@@
Assume there exist $w^{\star}\in\mathbb{R}^{D}$ and $\gamma>0$ with $\left\lVert w^{\star}\right\rVert=1$ such that
@@
Assume also there exists $R>0$ such that $\left\lVert x\right\rVert\leq R$ for all $(x,y)\in T$.
II-B Two-sided staleness and server-enforced staleness profile
The server maintains $w_{t}\in\mathbb{R}^{D}$ with initialization $w_{0}=0$ and the convention $w_{s}\equiv 0$ for all integers $s<0$.
Arrival (participation) sets
In each server round $t$, the server aggregates the set of client updates that arrive and are applied at round $t$. We denote this (possibly history-dependent) arrival set by $\mathcal{A}_{t}\subseteq\{1,\ldots,m\}$. No probabilistic model is imposed on the process $\{\mathcal{A}_{t}\}_{t\geq 0}$.
@@ II-B Two-sided staleness and server-enforced staleness profile
where $\delta_{i,t}$ is an effective downlink noise vector. Client $i$ runs the perceptron (any order; any finite number of passes), performs $k_{i,t}$ mistake updates, and outputs a local model $w_{i,t}$.
@@
The staleness-profile rule requires only that (i) the server can identify each realized total staleness $s_{i,t}$ (e.g., clients report the global model version used, or the server tags broadcasts), and (ii) the server stores the last $\tau$ iterates $\{w_{t-s}\}_{s=0}^{\tau}$.
@@
7:   Server learns/infers $(s^{\mathrm{dl}}_{i,t},s^{\mathrm{ul}}_{i,t})$ and sets $s_{i,t}\leftarrow s^{\mathrm{dl}}_{i,t}+s^{\mathrm{ul}}_{i,t}$.
8:   Client $i$ starts from $\tilde{w}^{\mathrm{dl}}_{i,t}=w_{t-s_{i,t}}+\delta_{i,t}$ and runs local perceptron updates on $T_{i}$, producing $(w_{i,t},k_{i,t})$.
@@
II-C Noise model
We model communication impairments via an effective additive noise abstraction with bounded conditional second moments [10]. For each round $t$, let $H_{t}$ denote the algorithmic history up to the start of round $t$, and let $F_{t}:=(H_{t},A_{t},\{s_{i,t}\}_{i\in A_{t}})$ denote the information available once the arrival set and stalenesses in round $t$ are revealed.
@@ III-B Main theorem
$\mathbb{E}[K_{A}]$ $\leq\frac{SR^{2}+\sqrt{S^{2}R^{4}+4\gamma^{2}SAV}}{2\gamma^{2}},$ (15)
$\leq\frac{SR^{2}}{\gamma^{2}}+\frac{\sqrt{SAV}}{\gamma}.$ (16)
Consequently,
$\frac{\mathbb{E}[K_{A}]}{A}\leq\frac{SR^{2}}{\gamma^{2}}\cdot\frac{1}{A}+\frac{\sqrt{SV}}{\gamma}\cdot\frac{1}{\sqrt{A}}\;\xrightarrow[A\to\infty]{}\;0.$ (17)
@@ IV-B Global recursions in expectation
Fix a round $t$. During round $t$ a set $A_{t}$ of updates arrives, with realized stalenesses $\{s_{i,t}\}_{i\in A_{t}}$ and buckets $\{B_{s,t}\}_{s=0}^{\tau}$.
@@
Condition on $\mathcal{F}_{t}$. Under this conditioning, the weights are fixed (they depend only on buckets) and the noise terms have zero conditional mean given $\mathcal{F}_{t}$, hence their conditional expectations vanish. Finally, by construction of the profile, for any scalar sequence $\{Z_{t-s}\}_{s=0}^{\tau}$,
@@
$\begin{split}\mathbb{E}[b_{t+1}\mid\mathcal{F}_{t}]&\leq\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}b_{t-s_{i,t}}+\sum_{s:\,\mathcal{B}_{s,t}=\emptyset}\pi_{s,t}b_{t-s}\\ &\quad+R^{2}\,\mathbb{E}[\kappa_{t}\mid\mathcal{F}_{t}]+V,\end{split}$ (*)
@@ IV-C Lyapunov potential and conclusion
Derivation of (28). Using Lemma 2 and the definition of $\Phi_{t}$,
$\Phi_{t+1}$ $=\sum_{j=0}^{\tau}c_{j}\,\mathbb{E}[a_{t+1-j}]=\mathbb{E}[a_{t+1}]+\sum_{j=1}^{\tau}c_{j}\,\mathbb{E}[a_{t+1-j}]$
$\geq\sum_{s=0}^{\tau}\alpha_{s}\,\mathbb{E}[a_{t-s}]+\gamma\,\mathbb{E}[\kappa_{t}]\;+\;\sum_{j=1}^{\tau}c_{j}\,\mathbb{E}[a_{t+1-j}]$
$=\sum_{s=0}^{\tau}\alpha_{s}\,\mathbb{E}[a_{t-s}]+\gamma\,\mathbb{E}[\kappa_{t}]\;+\;\sum_{s=0}^{\tau-1}c_{s+1}\,\mathbb{E}[a_{t-s}]$
$=\gamma\,\mathbb{E}[\kappa_{t}]+\sum_{s=0}^{\tau-1}(\alpha_{s}+c_{s+1})\,\mathbb{E}[a_{t-s}]+\alpha_{\tau}\mathbb{E}[a_{t-\tau}]$
$=\gamma\,\mathbb{E}[\kappa_{t}]+\sum_{s=0}^{\tau}c_{s}\,\mathbb{E}[a_{t-s}]=\Phi_{t}+\gamma\,\mathbb{E}[\kappa_{t}],$
where we used $\alpha_{s}+c_{s+1}=c_{s}$ for $s=0,\ldots,\tau-1$ and $c_{\tau}=\alpha_{\tau}$. Similarly, using Lemma 2 and the definition of $\Psi_{t}$,
$\Psi_{t+1}$ $=\sum_{j=0}^{\tau}c_{j}\,\mathbb{E}[b_{t+1-j}]=\mathbb{E}[b_{t+1}]+\sum_{j=1}^{\tau}c_{j}\,\mathbb{E}[b_{t+1-j}]$
$\leq\sum_{s=0}^{\tau}\alpha_{s}\,\mathbb{E}[b_{t-s}]+R^{2}\,\mathbb{E}[\kappa_{t}]+V\;+\;\sum_{j=1}^{\tau}c_{j}\,\mathbb{E}[b_{t+1-j}]$
$=\sum_{s=0}^{\tau}\alpha_{s}\,\mathbb{E}[b_{t-s}]+R^{2}\,\mathbb{E}[\kappa_{t}]+V\;+\;\sum_{s=0}^{\tau-1}c_{s+1}\,\mathbb{E}[b_{t-s}]$
$=R^{2}\,\mathbb{E}[\kappa_{t}]+V+\sum_{s=0}^{\tau-1}(\alpha_{s}+c_{s+1})\,\mathbb{E}[b_{t-s}]+\alpha_{\tau}\mathbb{E}[b_{t-\tau}]$
$=R^{2}\,\mathbb{E}[\kappa_{t}]+V+\sum_{s=0}^{\tau}c_{s}\,\mathbb{E}[b_{t-s}]=\Psi_{t}+R^{2}\,\mathbb{E}[\kappa_{t}]+V.$
@@
$=\mathbb{E}\!\left[\left\lVert\sum_{j=0}^{\tau}c_{j}w_{A-j}\right\rVert^{2}\right]$
@@
$=\left(\sum_{j=0}^{\tau}c_{j}\right)\,\sum_{j=0}^{\tau}c_{j}\,\mathbb{E}[\left\lVert w_{A-j}\right\rVert^{2}]$
@@
Thus $\Phi_{A}^{2}\leq S\Psi_{A}$. Combining with (29) gives $\gamma^{2}\mathbb{E}[K_{A}]^{2}\leq S(R^{2}\mathbb{E}[K_{A}]+AV)$; solving for $\mathbb{E}[K_{A}]$ yields (15)–(16).
@@ V Finite-Round Stabilization in the Noiseless Case
    $\mathbb{P}\big(i\in\mathcal{A}_{t},\ s_{i,t}=0\ \big|\ H_{t}\big)\ \geq\ p_{\min},$
(Fresh-mistake weight) The server uses any within-bucket rule (in the noiseless setting this may depend on the reported mistake counts $\{k_{i,t}\}$) such that whenever some fresh client makes a mistake in round $t$,
@@
$\mathbb{E}[T_{\mathrm{hit}}]\leq\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}},\qquad\mathbb{E}[T_{\mathrm{stab}}]\leq(\tau+1)\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}}.$
@@
If $\mathcal{E}_{t}$ does not occur then $\mathbf{1}\{\mathcal{E}_{t}\}=0$ and the inequality below is trivial. Otherwise, since $T=\biguplus_{i=1}^{m}T_{i}$, there exists at least one client index
$i_{t}\in\{1,\ldots,m\}\quad\text{such that}\quad\exists(x,y)\in T_{i_{t}}\ \text{with}\ y\left\langle w_{t},\,x\right\rangle\leq 0.$
@@
By (L1), $\mathbb{P}(G_{t}\mid H_{t})\geq p_{\min}$. On $G_{t}$, client $i_{t}$ starts from $w_{t}$ (noiseless downlink) and executes at least one complete pass over $T_{i_{t}}$. Because $w_{t}$ misclassifies at least one example in $T_{i_{t}}$, the first pass must contain at least one perceptron mistake update, hence $k_{i_{t},t}\geq 1$. Therefore, on $G_{t}$ there exists a fresh client that makes a mistake in round $t$, and by the fresh-mistake weight condition we have
@@
$\mathbb{E}[\kappa_{t}\mid H_{t}]\geq\alpha_{0}\,\mathbb{P}(G_{t}\mid H_{t})\geq\alpha_{0}p_{\min}\,\mathbf{1}\{\mathcal{E}_{t}\}.$
@@
$\sum_{t=0}^{\infty}\mathbb{P}(E_{t})\leq\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}}.$
@@
$\mathbb{E}[T_{\mathrm{hit}}]\leq\sum_{t=0}^{\infty}\mathbb{P}(E_{t})\leq\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}}.$
@@
$T_{\mathrm{stab}}=\sum_{t=0}^{\infty}\mathbf{1}\{t<T_{\mathrm{stab}}\}\leq\sum_{t=0}^{\infty}\sum_{j=0}^{\tau}\mathbf{1}\{\mathcal{E}_{t-j}\}=(\tau+1)\sum_{t=0}^{\infty}\mathbf{1}\{\mathcal{E}_{t}\}.$
@@ VII Discussion
Algorithm 1 implements a deterministic age-mixing rule: in every round the server forms a convex combination whose total-staleness distribution matches a chosen profile $\bm{\alpha}$, padding missing buckets with cached iterates. This removes the need to model delays or participation and explains Theorem 1: asynchrony enters only through the enforced mean staleness $\bar{s}$ (via $S=1+\bar{s}$), while link noise contributes separately through $V$ in (16).
A concrete next step is profile design: choose (or adapt) $\bm{\alpha}$ using online estimates of bucket occupancy to trade responsiveness (small $\bar{s}$) against feasibility (limited padding) while directly minimizing the bound (16). Moreover, Theorem 2 indicates that stabilization requires vanishing noise, so for persistent $V>0$ it is natural to couple aggregation with lightweight link protection (e.g., coding/retransmissions/denoising) that reduces the effective $V$ rather than expecting exact stabilization.

### v2

@@ Abstract
We study a semi-asynchronous client-server perceptron trained via iterative parameter mixing (IPM-style averaging): clients run local perceptron updates and a server forms a global model by aggregating the updates that arrive in each communication round. The setting captures three system effects in distributed deployments: (i) stale updates due to delayed model delivery and delayed application of client computations (two-sided version lag), (ii) partial participation (intermittent client availability), and (iii) imperfect communication on both downlink and uplink, modeled as effective zero-mean additive noise with bounded second moment. We introduce a server-side aggregation rule called staleness-bucket aggregation with padding that deterministically enforces a prescribed staleness profile over update ages without assuming any stochastic model for delays or participation. Under margin separability and bounded data radius, we prove a finite-horizon expected bound on the cumulative weighted number of perceptron mistakes over a given number of server rounds: the impact of delay appears only through the mean enforced staleness, whereas communication noise contributes an additional term that grows on the order of the square root of the horizon with the total noise energy. In the noiseless case, we show how a finite expected mistake budget yields an explicit finite-round stabilization bound under a mild fresh-participation condition. Synthetic experiments match the predicted square root of horizon noise scaling.
@@ I Introduction
Practical distributed deployments must contend with three persistent phenomena. These effects are ubiquitous and can break classical perceptron guarantees, motivating a delay/noise-aware mistake bound handling: (i) staleness (version lag), as in bounded-staleness frameworks such as SSP [6, 7] and online learning with delayed feedback [14, 15]; (ii) partial participation, where only a subset of clients return updates each round [5, 11, 12]; and (iii) communication noise on downlink/uplink, especially in wireless settings [8, 9, 10].
@@
II System Model and Staleness-Profile IPM Perceptron
@@
The dataset $T$ is $\gamma$-margin separable with radius $R$ if there exist $w^{\star}\in\mathbb{R}^{D}$ with $\left\lVert w^{\star}\right\rVert=1$ and constants $\gamma>0$, $R>0$ such that
@@
with $\left\lVert x\right\rVert\leq R$ for all $(x,y)\in T$.
II-A Two-sided staleness and server-enforced staleness profile
The server maintains $w_{t}\in\mathbb{R}^{D}$ with initialization $w_{0}=0$ and the convention $w_{s}\equiv 0$ for all integers $s<0$. In each server round $t$, the server aggregates the set of client updates that arrive and are applied at round $t$. We denote this (possibly history-dependent) arrival set by $\mathcal{A}_{t}\subseteq\{1,\ldots,m\}$. No probabilistic model is imposed on the process $\{\mathcal{A}_{t}\}_{t\geq 0}$.
@@ II-A Two-sided staleness and server-enforced staleness profile
where $\delta_{i,t}$ is an effective downlink noise vector. Client $i$ runs the perceptron (any order; for at least one epoch), performs $k_{i,t}$ mistake updates, and outputs a local model $w_{i,t}$.
@@
The staleness-profile rule requires only that (i) the server can identify each realized total staleness $s_{i,t}$ (e.g., clients report the global model version used, or the server tags broadcasts), and (ii) the server stores the last $\tau+1$ iterates $\{w_{t-s}\}_{s=0}^{\tau}$.
@@
7:   Server learns/infers $(s^{\mathrm{dl}}_{i,t},s^{\mathrm{ul}}_{i,t})$; set $a\leftarrow t-s^{\mathrm{ul}}_{i,t}$ and $s_{i,t}\leftarrow s^{\mathrm{dl}}_{i,t}+s^{\mathrm{ul}}_{i,t}$.
8:   Client $i$ started local computation in round $a$ from $\tilde{w}^{\mathrm{dl}}_{i,t}=w_{a-s^{\mathrm{dl}}_{i,t}}+\delta_{i,t}$ and runs local perceptron on $T_{i}$ for at least one epoch, producing $(w_{i,t},k_{i,t})$.
@@
II-B Noise model
We model communication impairments via an effective additive noise abstraction with bounded conditional second moments [10]. For each round $t$, let $H_{t}$ denote the algorithmic history up to the start of round $t$, and let $\mathcal{F}_{t}:=(H_{t},\mathcal{A}_{t},\{s_{i,t}\}_{i\in\mathcal{A}_{t}})$ denote the information available once the arrival set and stalenesses in round $t$ are revealed.
@@ III-B Main theorem
$\mathbb{E}[K_{A}]\leq\frac{SR^{2}}{\gamma^{2}}+\frac{\sqrt{SAV}}{\gamma}.$ (15)
@@
As a corollary, (15) implies $\mathbb{E}[K_{A}]/A\to 0$ as $A\to\infty$, i.e., a vanishing expected per-round weighted mistake rate.
@@ IV-B Global recursions in expectation
Fix a round $t$. During round $t$ a set $\mathcal{A}_{t}$ of updates arrives, with realized stalenesses $\{s_{i,t}\}_{i\in\mathcal{A}_{t}}$ and buckets $\{\mathcal{B}_{s,t}\}_{s=0}^{\tau}$.
@@
Condition on $\mathcal{F}_{t}$. Under this conditioning, the weights are fixed (they depend only on buckets) and the noise terms have zero conditional mean given $\mathcal{F}_{t}$ (for $\xi_{i,t}$ this follows from (10) by iterated expectation over $w_{i,t}$), hence their conditional expectations vanish. Finally, by construction of the profile, for any scalar sequence $\{Z_{t-s}\}_{s=0}^{\tau}$,
@@
$\begin{split}\mathbb{E}[b_{t+1}\mid\mathcal{F}_{t}]&\leq\sum_{i\in\mathcal{A}_{t}}\mu_{i,t}b_{t-s_{i,t}}+\sum_{s:\,\mathcal{B}_{s,t}=\emptyset}\pi_{s,t}b_{t-s}\\ &\quad+R^{2}\,\mathbb{E}[\kappa_{t}\mid\mathcal{F}_{t}]+V,\end{split}$
@@ IV-C Lyapunov potential and conclusion
Using $c_{j}-c_{j+1}=\alpha_{j}$ (with $c_{\tau+1}:=0$), one can rewrite $\Phi_{t+1}=\Phi_{t}+\mathbb{E}[a_{t+1}]-\sum_{s=0}^{\tau}\alpha_{s}\mathbb{E}[a_{t-s}]$ (and similarly for $\Psi_{t}$); substituting Lemma 2 yields the one-step bounds
@@
Thus $\Phi_{A}^{2}\leq S\Psi_{A}$. Combining with (27) gives $\gamma^{2}\mathbb{E}[K_{A}]^{2}\leq S(R^{2}\mathbb{E}[K_{A}]+AV)$; solving for $\mathbb{E}[K_{A}]$ yields (15).
@@ V Finite-Round Stabilization in the Noiseless Case
    $\mathbb{P}\big(i\in\mathcal{A}_{t},\ s_{i,t}=0\ \big|\ H_{t}\big)\ \geq\ p_{\min}$
The server uses any within-bucket rule (in the noiseless setting this may depend on the reported mistake counts $\{k_{i,t}\}$) such that whenever some fresh client makes a mistake in round $t$,
@@
$\mathbb{E}[T_{\mathrm{hit}}]$ $\leq\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}},$ (31)
$\mathbb{E}[T_{\mathrm{stab}}]$ $\leq(\tau+1)\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}}.$
@@
If $\mathcal{E}_{t}$ does not occur then $\mathbf{1}\{\mathcal{E}_{t}\}=0$ and the inequality below is trivial. Otherwise, since $T=\biguplus_{i=1}^{m}T_{i}$, there exists a client index $i_{t}\in\{1,\ldots,m\}$ and some $(x,y)\in T_{i_{t}}$ such that $y\left\langle w_{t},\,x\right\rangle\leq 0$.
@@
By (L1), $\mathbb{P}(G_{t}\mid H_{t})\geq p_{\min}$. On $G_{t}$, client $i_{t}$ starts from $w_{t}$ (noiseless downlink) and executes at least one epoch over $T_{i_{t}}$. Because $w_{t}$ misclassifies at least one example in $T_{i_{t}}$, the first epoch must contain at least one perceptron mistake update, hence $k_{i_{t},t}\geq 1$. Therefore, on $G_{t}$ there exists a fresh client that makes a mistake in round $t$, and by the fresh-mistake weight condition we have
@@
$\mathbb{E}[\kappa_{t}\mid H_{t}]\geq\alpha_{0}p_{\min}\,\mathbf{1}\{\mathcal{E}_{t}\}.$
@@
$\sum_{t=0}^{\infty}\mathbb{P}(\mathcal{E}_{t})\leq\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}}.$
@@
$\mathbb{E}[T_{\mathrm{hit}}]\leq\sum_{t=0}^{\infty}\mathbb{P}(\mathcal{E}_{t})\leq\frac{SR^{2}}{\alpha_{0}p_{\min}\gamma^{2}}.$
@@
$T_{\mathrm{stab}}$ $=\sum_{t=0}^{\infty}\mathbf{1}\{t<T_{\mathrm{stab}}\}$
$\leq\sum_{t=0}^{\infty}\sum_{j=0}^{\tau}\mathbf{1}\{\mathcal{E}_{t-j}\}$
$=(\tau+1)\sum_{t=0}^{\infty}\mathbf{1}\{\mathcal{E}_{t}\}.$
@@ VII Discussion
Algorithm 1 implements a deterministic age-mixing rule: in every round the server forms a convex combination whose total-staleness distribution matches a chosen profile $\bm{\alpha}$, padding missing buckets with cached iterates. This removes the need to model delays or participation and explains Theorem 1: asynchrony enters only through the enforced mean staleness $\bar{s}$ (via $S=1+\bar{s}$), while link noise contributes through $V$ in (15). Since $S$ appears linearly in $SR^{2}/\gamma^{2}$ and inside the square root in $\sqrt{SAV}/\gamma$, reducing $\bar{s}$ improves both the noiseless plateau and the noisy $\sqrt{A}$ slope.
A practical profile-design heuristic is to estimate bucket-occupancy frequencies online and choose $\bm{\alpha}$ supported on stalenesses that are reliably present; otherwise that mass is frequently assigned to cached iterates, which is safe but reduces the fraction of new information mixed each round. Among feasible profiles, (15) suggests minimizing $\bar{s}$ to shrink both terms. Finally, (15) also reveals a noise-limited regime: once $\sqrt{SAV}/\gamma$ dominates $SR^{2}/\gamma^{2}$, additional rounds yield diminishing returns unless the effective $V$ is reduced.